# Mixing Individual and Collective Behaviours to Predict Out-of-Routine Mobility

### v1

@@ I Introduction
Both deep learning and Markov models are trained on individual mobility trajectories, which are typically inherently predictable as people tend to visit previously visited locations at regular times [28, 29, 30, 31, 32, 33, 19]. However, in some cases, individuals may have a marked preference for exploring new destinations [31, 32, 34, 35] or be forced to alter their routine due to external factors such as job loss, health issues, natural disasters, or epidemics [36, 37, 38, 39, 40, 41, 42, 43, 44, 9]. Predicting such out-of-routine mobility is a challenge for statistical models because, being designed to capture regular patterns in individual trajectories, they often memorise training data rather than learning generalised mobility behaviours [17, 45]. A large body of literature on human behaviour across various contexts, such as social networks [46, 47, 48, 49], financial networks [50, 51, 52], and voting and political polarization [53, 54, 55], indicates that an individual’s decisions are significantly influenced by the behaviour of the group or community they are exposed to [56, 57, 58]. This suggests that information about collective behaviours also holds predictive power for individuals. However, the potential of combining individual and collective behaviours to enhance human mobility prediction remains largely unexplored.
@@ Results
A spatio-temporal point is a pair $p=(i,t)$, where $i$ represents a geographic location and $t$ the time of the visit. We define a trajectory $P=\{p_{1},p_{2},\dots,p_{n}\}$ as a daily time-ordered sequence of $n$ spatio-temporal points. Each individual user $u$ has a set of $N$ historical trajectories $\mathcal{H}^{(u)}=\{P_{1},\dots,P_{N}\}$ from which we compute $I^{(u)}_{i}$, representing the set of transition probabilities of user $u$ starting from location $i$ (see Figure 1A and Methods for details). By aggregating the trajectories of all individuals, we calculate the collective origin-destination matrix $C$ (see Figure 1B), with $C_{i}$ representing the probability distribution of all transitions made by any individual starting from location $i$.
@@
where $l_{1},l_{2},\dots l_{n}$ are the transition probabilities of $M_{i}^{(u)}$ in Equation 2. When a transition from location $i$ is not represented in $\mathcal{H}^{(u)}$, the probability distribution $I^{(u)}_{i}$ is empty, and we set $S^{(u)}_{i}=1$ to indicate maximum uncertainty. In such instances, no historical data is available for location $i$, and the model prediction relies solely on collective information in $C_{i}$. Figure 1C-E illustrates how model $M_{i}^{(u)}$ works.
@@
We employ the notations $I$, $C$, and $M$ to represent the models exploiting individual, collective, and their combination. To derive $I$ and $C$ for real individuals, we use privacy-enhanced GPS trajectories collected in Boston, Seattle, and New York City (NYC) from January 3rd to March 1st, 2020 (see Supplementary Note 1). We tessellate the cities into rectangular tiles (locations) of $1.2$ km $\times$ $609.4$ m, corresponding to GeoHashes of level 6. These tiles are the basis for mapping user’s stops and computing users’ trajectories (see Methods). To address potential under- and over-representation issues, we filter out trajectories with less than four points ($|P^{(u)}|<4$), users with fewer than two trajectories ($|\mathcal{H}^{(u)}|<2$), and remove the top 95th percentile of the most represented users (see Supplementary Note 1). Figure 1A-B shows examples of $I$ and $C$ derived from the GPS dataset for the city of Boston.
@@
Table 1 and Figure 2A show that $M$ significantly enhances accuracy compared to $I$ and $C$ and obtains comparable performances with deep learning baselines like Recurrent Neural Networks (RNNs) [61, 17]. $M$ exhibits relative improvements in ACC@5 over $I$ of +15% in NYC, +13% in Seattle, and +12% in Boston. Furthermore, $M$ outperforms $C$ with improvements of +35% in NYC, +31% in Seattle, and +21% in Boston. Despite relying on a large number of parameters, RNNs exhibit only marginal relative improvements over $M$: +2% in NYC, +2% in Seattle, and +1% in Boston (see Supplementary Note 6). Therefore, the dynamic interplay between individual and collective information enables model $M$ to significantly enhance the overall predictive performance compared to models relying exclusively on individual ($I$) or collective ($C$) information.
@@
We quantify trajectory train-test overlap with the Longest Common Sub-Trajectory (LCST) [45], which evaluates shared sub-sequences between two trajectories, considering the order and frequency of visits. LCST ranges between 0 and 1 and measures how much of a test trajectory the model has already observed in the training set. A high LCST indicates the presence of recurrent mobility patterns in an individual’s trajectories, while a low LCST indicates instances of an individual’s out-of-routine mobility behaviours. We stratify test trajectories into bins based on LCST, spanning overlap ranges of 0-20%, 20-40%, 40-60%, 60-80%, and 80-100% (see Methods for details). For example, a test trajectory falls within the 0-20% bin when its maximum overlap with any training trajectory has an LCST $\in[0,0.2]$. The distribution of test trajectories across the bins is not uniform, as bins 0-20% and 80-100% contain fewer trajectories than bins 20-80% (see Supplementary Note 1).
Table 1 and Figure 2B provide models’ performance within each trajectory overlap bin. $M$ significantly outperforms $I$ and $C$ for intermediate levels of overlap, with improvements across cities up to +16% (20-40% overlap), +13% (40-60% overlap), and around $+2$% (60-80% overlap). $C$ surpasses $M$ for 0-20% overlap only, with improvements up to 16% across the cities, while $I$ outperforms $M$ by only around 1% for 80-100% overlap. Figure 2C shows that the median value of $1-S^{(u)}_{i}$ distribution increases with trajectory overlap, i.e., the higher the trajectory overlap, the higher the confidence of model $M$ on $I$ and, consequently, the lower the reliance on $C$. In other words, when historical information about the individual is unreliable due to significant differences from the current trajectory (low trajectory overlap), using collective information is crucial for making reasonable predictions. When the current trajectory closely resembles those in the training set (high trajectory overlap), it is best to rely on individual information. Note that RNNs achieve a low accuracy on low trajectory overlap bins and high accuracy on high overlap bins (Figure 2B), confirming that deep learning models struggle to predict out-of-routine mobility. To verify the robustness of our findings, we carry out the same analysis using more fine-grained tessellations, obtaining comparable results (see Supplementary Note 6).
@@
To investigate potential spatial dependencies in the accuracy of $I$, $C$ and $M$, we compute ACC@5 specifically for the subset of test transitions originating from a location $i$. We refer to it as ACC@5i. In Figure 3A, we present the spatial distributions of ACC@5i for the city of Boston. We find similar distributions for Seattle and NYC (see Supplementary Note 2).
In scenarios with low trajectory overlap (0-40%), the ACC@5i distributions for $M$ and $C$ are consistently shifted to higher values compared to $I$ (see Figure 3A). While model $C$ performs best for trajectories with 0-20% overlap, model $M$ shows the highest performance in the 20-40% overlap range, benefiting from collective information from $C$. As trajectory overlap increases, the ACC@5i distributions for $M$ and $I$ become more similar, while the distribution for $C$ widens and remains skewed towards lower accuracy (see Figure 3A).
Figure 3C shows that when out-of-routine is predominant (0-40% overlap), there is considerable spatial heterogeneity in the distribution of ACC@5i across different locations. This is particularly pronounced when using collective information for predictions, as in $C$ and $M$. We quantify this spatial property using Moran’s index [63, 64] and find that models $C$ and $M$ exhibit a significant and positive spatial autocorrelation, indicating that locations where the models are accurate are spatially close (see Figure 3B-C and Supplementary Figure S4-S5).
In particular, we observe that locations with high accuracies for $C$ and $M$ are clustered in the proximity of critical urban areas. For instance, in Boston around downtown and Logan International Airport (see Figure 3C), in NYC in Manhattan, and in Seattle around downtown (see Supplementary Figures S4-S5).
@@
where $L^{(C)}$ is the set of unique locations in $C$. $S_{i}^{(C)}$ is high when locations people visited from $i$ have similar visitation probabilities, indicating a diverse range of destinations; $S_{i}^{(C)}$ is low when people predominantly visit one location from $i$, indicating a marked collective preference for a specific destination. $S^{(C)}_{i}$ is strongly negatively correlated with ACC@5i (Pearson correlation of $\rho=-0.85$), indicating that locations from which $C$ provides the most accurate next location predictions are those with the lowest entropy $S^{(C)}_{i}$ (see Figure 4A and Supplementary Figures S8-S9). Notably, we find that locations with low $S_{i}^{(C)}$ are clustered in proximity to specific urban areas, which we hypothesise to be locations hosting key commercial, financial, and cultural venues.
To verify this hypothesis, we first collect from OpenStreetMap [65] the number of points of interest (POIs) in each location $i$, $W_{i}$. We then split each city into two areas: one comprising locations within a geographical distance $D$ from the location $i^{*}$ with the maximum number of POIs ($W_{i}^{*}$), and the other consisting of locations farther away from $i^{*}$ by a distance greater than $D$ (see Figure 4B for $D=2$ km in Boston).
We find that the distribution of $S_{i}^{(C)}$ skews towards lower values when location $i$ is within distance $D$ to $i^{*}$ (orange area in Figure 4B). Additionally, individuals’ movements that originate within this distance tend to travel shorter distances than those originating farther away (see Figure 4C). These results support our hypothesis that mobility in dense areas of POIs is less spatially dispersed, with the probability of travelling to a destination concentrated in a small subset of locations. We see similar results in Seattle and NYC (Supplementary Figures S6-7). For statistical robustness, we verify that these results are not a consequence of possible biases, such as areas close to POIs having a larger sample size of transitions in the dataset (see Supplementary Note 4 and Supplementary Figures S10-14).
@@ Methods
We download from OpenStreetMap (OSM) data about points of interest (POIs) in New York City, Seattle, and Boston. POIs describe public venues such as restaurants, and parks in a city. We employ a dictionary of amenities covering a large set of public venues (see Supplementary Note 1 for the specific entries used). We compute the number of POIs extracted from OSM in each GeoHash tile $i$ and refer to it as $W_{i}$. The maps of POIs for Boston, Seattle and New York City can be seen in Supplementary Figure S6.
@@
We normalise the LCST score within the range $[0,1]$, and we assign each trajectory to one of the following five bins: 0-20%, 20-40%, 40-60%, 60-80%, and 80-100%, based on their LCST score [45]. For instance, a test trajectory with an LCST score less than or equal to $0.2$ will be assigned to the bin of trajectories with 0-20% LCST. We exclude trajectories with an exact 0% overlap as they are primarily consisting of individuals remaining in the same location, for a small set of locations. Refer to Supplementary Note 1 for further details. The distribution of LCST for test trajectories is detailed in Supplementary Note 1.
Reporting summary
Further information on research design is available in the Nature Portfolio Reporting Summary linked to this article.

### v2

@@ Abstract
Significance Statement
Crowds significantly influence individual decisions, as evidenced by studies on collective intelligence, social psychology, and behavioural economics. Existing models that predict individual patterns struggle to predict unexpected behaviours as they tend to memorise individual preferences and lack generalisation capabilities. Our approach mixes collective and individual decisions to overcome these limitations while offering interpretability and stronger performance in out-of-routine mobility prediction. We find that collective information is crucial in forecasting these unconventional choices by capturing broader patterns and trends that individual-based models overlook. In addition, we find that collective information makes the model robust to external shocks and reveals that out-of-routine mobility patterns close to areas with a higher density of points of interest are more predictable and more localised.
@@ I Introduction
Both deep learning and Markov models are trained on individual mobility trajectories, which are typically inherently predictable as people tend to visit previously visited locations at regular times [28, 29, 30, 31, 32, 33, 19]. However, in some cases, individuals may have a marked preference for exploring new destinations [31, 32, 34, 35, 36] or be forced to alter their routine due to external factors such as job loss, health issues, natural disasters, or epidemics [37, 38, 39, 40, 41, 42, 43, 44, 45, 9]. Predicting such out-of-routine mobility is a challenge for statistical models because, being designed to capture regular patterns in individual trajectories, they often memorise training data rather than learning generalised mobility behaviours [17, 46]. A large body of literature on human behaviour across various contexts, such as social networks [47, 48, 49, 50], financial networks [51, 52, 53], and voting and political polarization [54, 55, 56], indicates that an individual’s decisions are significantly influenced by the behaviour of the group or community they are exposed to [57, 58, 59]. This suggests that information about collective behaviours also holds predictive power for individuals. However, the potential of combining individual and collective behaviours to enhance human mobility prediction remains largely unexplored.
@@ Results
A spatio-temporal point is a pair $p=(i,t)$, where $i$ represents a geographic location and $t$ the time of the visit. We define a trajectory $P=\{p_{1},p_{2},\dots,p_{n}\}$ as a daily time-ordered sequence of $n$ spatio-temporal points. Each individual user $u$ has a set of $N$ historical trajectories $\mathcal{H}^{(u)}=\{P_{1},\dots,P_{N}\}$ from which we compute $I^{(u)}_{i}$, representing the set of transition probabilities of user $u$ starting from location $i$ (see Fig. 1A and Methods for details). By aggregating the trajectories of all individuals, we calculate the collective origin-destination matrix $C$ (see Fig. 1B), with $C_{i}$ representing the probability distribution of all transitions made by any individual starting from location $i$.
@@
where $l_{1},l_{2},\dots l_{n}$ are the transition probabilities of $M_{i}^{(u)}$ in Equation 2. When a transition from location $i$ is not represented in $\mathcal{H}^{(u)}$, the probability distribution $I^{(u)}_{i}$ is empty, and we set $S^{(u)}_{i}=1$ to indicate maximum uncertainty. In such instances, no historical data is available for location $i$, and the model prediction relies solely on collective information in $C_{i}$. Fig. 1C-E illustrates how model $M_{i}^{(u)}$ works.
@@
We employ the notations $I$, $C$, and $M$ to represent the models exploiting individual, collective, and their combination. To derive $I$ and $C$ for real individuals, we use privacy-enhanced GPS trajectories collected in Boston, Seattle, and New York City (NYC) from January 3rd to March 1st, 2020 (see Supplementary Note S1, Supplementary Table S1, Supplementary Figure S1). We tessellate the cities into rectangular tiles (locations) of $1.2$ km $\times$ $609.4$ m, corresponding to GeoHashes of level 6. These tiles are the basis for mapping user’s stops and computing users’ trajectories (see Methods). To address potential under- and over-representation issues, we filter out trajectories with less than four points ($|P^{(u)}|<4$), users with fewer than two trajectories ($|\mathcal{H}^{(u)}|<2$), and remove the top 95th percentile of the most represented users (see Supplementary Note S1, Supplementary Figure S3). Fig. 1A-B shows examples of $I$ and $C$ derived from the GPS dataset for the city of Boston.
@@
Table 1 and Fig. 2A show that $M$ significantly enhances accuracy compared to $I$ and $C$ and obtains comparable performances with deep learning baselines like Recurrent Neural Networks (RNNs) [62, 17]. $M$ exhibits relative improvements in ACC@5 over $I$ of +15% in NYC, +13% in Seattle, and +12% in Boston. Furthermore, $M$ outperforms $C$ with improvements of +35% in NYC, +31% in Seattle, and +21% in Boston. Despite relying on a large number of parameters, RNNs exhibit only marginal relative improvements over $M$: +2% in NYC, +2% in Seattle, and +1% in Boston (see Supplementary Note S6, Supplementary Table S3). Therefore, the dynamic interplay between individual and collective information enables model $M$ to significantly enhance the overall predictive performance compared to models relying exclusively on individual ($I$) or collective ($C$) information.
@@
We quantify trajectory train-test overlap with the Longest Common Sub-Trajectory (LCST) [46], which evaluates shared sub-sequences between two trajectories, considering the order and frequency of visits. LCST ranges between 0 and 1 and measures how much of a test trajectory the model has already observed in the training set. A high LCST indicates the presence of recurrent mobility patterns in an individual’s trajectories, while a low LCST indicates instances of an individual’s out-of-routine mobility behaviours. We stratify test trajectories into bins based on LCST, spanning overlap ranges of 0-20%, 20-40%, 40-60%, 60-80%, and 80-100% (see Methods for details). For example, a test trajectory falls within the 0-20% bin when its maximum overlap with any training trajectory has an LCST $\in[0,0.2]$. The distribution of test trajectories across the bins is not uniform, as bins 0-20% and 80-100% contain fewer trajectories than bins 20-80% (see Supplementary Note S1, Supplementary Figure S3).
Table 1 and Fig. 2B provide models’ performance within each trajectory overlap bin. $M$ significantly outperforms $I$ and $C$ for intermediate levels of overlap, with improvements across cities up to +16% (20-40% overlap), +13% (40-60% overlap), and around $+2$% (60-80% overlap). $C$ surpasses $M$ for 0-20% overlap only, with improvements up to 16% across the cities, while $I$ outperforms $M$ by only around 1% for 80-100% overlap. Fig. 2C shows that the median value of $1-S^{(u)}_{i}$ distribution increases with trajectory overlap, i.e., the higher the trajectory overlap, the higher the confidence of model $M$ on $I$ and, consequently, the lower the reliance on $C$. In other words, when historical information about the individual is unreliable due to significant differences from the current trajectory (low trajectory overlap), using collective information is crucial for making reasonable predictions. When the current trajectory closely resembles those in the training set (high trajectory overlap), it is best to rely on individual information. Note that RNNs achieve a low accuracy on low trajectory overlap bins and high accuracy on high overlap bins (Fig. 2B), confirming that deep learning models struggle to predict out-of-routine mobility. To verify the robustness of our findings, we carry out the same analysis using more fine-grained tessellations, obtaining comparable results (see Supplementary Note S6).
@@
To investigate potential spatial dependencies in the accuracy of $I$, $C$ and $M$, we compute ACC@5 specifically for the subset of test transitions originating from a location $i$. We refer to it as ACC@5i. In Fig. 3A, we present the spatial distributions of ACC@5i for the city of Boston. We find similar distributions for Seattle and NYC (see Supplementary Note S2, Supplementary Figure S4, Supplementary Figure S5).
In scenarios with low trajectory overlap (0-40%), the ACC@5i distributions for $M$ and $C$ are consistently shifted to higher values compared to $I$ (see Fig. 3A). While model $C$ performs best for trajectories with 0-20% overlap, model $M$ shows the highest performance in the 20-40% overlap range, benefiting from collective information from $C$. As trajectory overlap increases, the ACC@5i distributions for $M$ and $I$ become more similar, while the distribution for $C$ widens and remains skewed towards lower accuracy (see Fig. 3A).
Fig. 3C shows that when out-of-routine is predominant (0-40% overlap), there is considerable spatial heterogeneity in the distribution of ACC@5i across different locations. This is particularly pronounced when using collective information for predictions, as in $C$ and $M$. We quantify this spatial property using Moran’s index [64, 65] and find that models $C$ and $M$ exhibit a significant and positive spatial autocorrelation, indicating that locations where the models are accurate are spatially close (see Fig. 3B-C and Supplementary Fig. S4-S5).
In particular, we observe that locations with high accuracies for $C$ and $M$ are clustered in the proximity of critical urban areas. For instance, in Boston around downtown and Logan International Airport (see Fig. 3C), in NYC in Manhattan, and in Seattle around downtown (see Supplementary Fig. S4-S5).
@@
where $L^{(C)}$ is the set of unique locations in $C$. $S_{i}^{(C)}$ is high when locations people visited from $i$ have similar visitation probabilities, indicating a diverse range of destinations; $S_{i}^{(C)}$ is low when people predominantly visit one location from $i$, indicating a marked collective preference for a specific destination. $S^{(C)}_{i}$ is strongly negatively correlated with ACC@5i (Pearson correlation of $\rho=-0.85$), indicating that locations from which $C$ provides the most accurate next location predictions are those with the lowest entropy $S^{(C)}_{i}$ (see Fig. 4A and Supplementary Figure S8-S9). Notably, we find that locations with low $S_{i}^{(C)}$ are clustered in proximity to specific urban areas, which we hypothesise to be locations hosting key commercial, financial, and cultural venues.
To verify this hypothesis, we first collect from OpenStreetMap [66] the number of points of interest (POIs) in each location $i$, $W_{i}$. We then split each city into two areas: one comprising locations within a geographical distance $D$ from the location $i^{*}$ with the maximum number of POIs ($W_{i}^{*}$), and the other consisting of locations farther away from $i^{*}$ by a distance greater than $D$ (see Fig. 4B for $D=2$ km in Boston).
We find that the distribution of $S_{i}^{(C)}$ skews towards lower values when location $i$ is within distance $D$ to $i^{*}$ (orange area in Fig. 4B). Additionally, individuals’ movements that originate within this distance tend to travel shorter distances than those originating farther away (see Fig. 4C). These results support our hypothesis that mobility in dense areas of POIs is less spatially dispersed, with the probability of travelling to a destination concentrated in a small subset of locations. We see similar results in Seattle and NYC (Supplementary Figure S6-S7). For statistical robustness, we verify that these results are not a consequence of possible biases, such as areas close to POIs having a larger sample size of transitions in the dataset (see Supplementary Note S4 and Supplementary Figure S10-S14).
@@ Methods
We download from OpenStreetMap (OSM) data about points of interest (POIs) in New York City, Seattle, and Boston. POIs describe public venues such as restaurants, and parks in a city. We employ a dictionary of amenities covering a large set of public venues (see Supplementary Note S1 for the specific entries used). We compute the number of POIs extracted from OSM in each GeoHash tile $i$ and refer to it as $W_{i}$. The maps of POIs for Boston, Seattle and New York City can be seen in Supplementary Fig. S6.
@@
We normalise the LCST score within the range $[0,1]$, and we assign each trajectory to one of the following five bins: 0-20%, 20-40%, 40-60%, 60-80%, and 80-100%, based on their LCST score [46]. For instance, a test trajectory with an LCST score less than or equal to $0.2$ will be assigned to the bin of trajectories with 0-20% LCST. We exclude trajectories with an exact 0% overlap as they are primarily consisting of individuals remaining in the same location, for a small set of locations. Refer to Supplementary Note S1, Supplementary Figure S2, Supplementary Table S2, Supplementary Figure S3 for further details. The distribution of LCST for test trajectories is detailed in Supplementary Note S1 and Supplementary Figure S2.